OpenStudy (barrelracer17):

What is the distance between (6, 2) and (-3, -2)? 9 3 square root of 97 5

OpenStudy (barrelracer17):

@DaBest21

pooja195 (pooja195):

the distance formula \[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We are trying to figure out the distance between the 2 point right? This formula will help us find the distance understood?

OpenStudy (barrelracer17):

yeah

pooja195 (pooja195):

So you do know that a coordinate is set up like this right? (x,y) <---

OpenStudy (barrelracer17):

yes

pooja195 (pooja195):

Ok next we need know that this is how we determine what goes where.... (x1,y1)(x2,y2) understood?

OpenStudy (barrelracer17):

okay

pooja195 (pooja195):

\[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] x1=? x2=? y1=? y2=?

pooja195 (pooja195):

(6, 2) and (-3, -2) x1=6 x2=-3 y1=2 y2=-2 the distance formula \[\huge~\rm~\sqrt{(-3-6)^2+(-2-2)^2}\] im sure you can solve it from here.

OpenStudy (barrelracer17):

thank you poooja195

pooja195 (pooja195):

No problem :D